The tens digit of a two-digit number is 3 times the units digit. The diference of the digits is 6. What are the 2 equations?
Answer by rapaljer(4671)

Let t = tens digit
u = units digit

The two equations:
t=3u, (which by the way implies that t is larger than u!)
t-u= 6

You don't want to solve the problem? It's no extra charge!!

Substitute t=3u into the second equation:
t-u= 6
(____)-u = 6
3u - u = 6
2u = 6
u= 3

t=3u
t=3*3= 9

The number is 93!

R^2 at SCC